# Need Help? Designing Proactive AI Assistants for Programming

### v1

@@ 7.1.2 Expanding design considerations
(6) Allow users to decide when they want proactive assistance. Multiple participants across conditions recommended having an option for users to decide when to turn the proactive suggestions (e.g., “an opt-in feature” or “could change the frequency”), particularly because “sometimes I found it useful and sometimes I didn’t”. Since the goal of our study was to investigate the usage patterns and effects of a proactive assistant, providing the user the option to turn off the proactive assistant would not have been possible. However, this feature could be important for facilitating appropriate long-term usage, as a participant was concerned about potential over-reliance on proactive suggestions, which “can lead to a lack of understanding of a codebase. Therefore, I don’t want constant suggestions.”
@@
(7) Incorporate additional context when generating suggestions beyond user code. While many participants already appreciated the benefits of proactive suggestions, they had further ideas for improving the suggestion content by incorporating other aspects of user behavior. For example, participants suggested that the proactive assistant could take into account “the functions I was using’’, “what the user is focusing on, typing, or clicking,” or “recently updated code regions.” Another participant suggested that it may be beneficial for the user to incorporate further context by providing a natural language of their task—i.e., “if I can enter the task description somewhere as a constant part of the prompt [as] some suggestions have a slight misunderstanding of the problem”. These recommendations demonstrate that augmenting the existing code context with additional user information may be able to further tailor suggestions to user needs.
@@
(8) Flexibly varying the amount of information shown. Since the current proactive assistant provides one implementation option and a fixed number of suggestions each time, participants noted the possibility of scaling the amount of information up and down. For example, the assistant could provide more implementation options in the proactive suggestions, e.g., “multiple ways of solving a problem,” particularly when different factors may be important (e.g., time versus space efficiency). However, there may also be times when the assistant could scale down the number of suggestions shown (e.g., “if everything was good, then tell me that everything is good.”)
7.2 Limitations
@@
There are a few limitations to our study results and setup which we discuss here. First, our study is time-limited and consisted of tasks that were given to the participants rather than the tasks being self-motivated by the participants. The coding tasks we relied on do not span the entire set of tasks a professional programmer might encounter in their work. The tasks were also easily measured in the number of test cases passed, which may not be the best metric in practice. Second, the participants in our study were entirely comprised of undergraduate and graduate students, which is not fully representative of the audience that uses AI tools for programming. Third, our implementation of the proactive assistant was integrated into a web-based coding environment which does not contain the full features of an enterprise IDE. Fourth, the conditions that we considered in our controlled lab study were limited to chat-based systems, since not all participants had experience using GitHub Copilot.
@@
Further work is necessary to understand the nuances of proactive assistants on a wider range of user backgrounds and coding tasks. Future work could consider evaluating the proactive chat assistants in the presence of other programming tools like code completion to understand the interplay between different tools and their relative effects on productivity.
@@ A.1 Instructions
  • You can run your code by pressing the "Run" button and the output will be in the output box at the bottom in grey.
@@
  • Pressing the "Clear" button on the top of the chat window will clear the current chat history. Chat history will be automatically cleared when you move to the next task.
@@
  • Previewing: Click "Preview" to see how the proactive assistant incorporates the suggestion into your code, which can then accept or hide the changes. Use the copy icon to copy the code to your clipboard.
@@
You can review this tutorial again by clicking on the “Show Instructions” button at the top of the page, then clicking "Next".
A.2 Post-study Questionnaire
@@ B.1 Tasks
You are an application developer working on the backend of a to-do list app. You are working with some existing code your colleague left off. Each task takes in a ‘description‘ (string), ‘category‘ (string), ‘priority‘ (string, ‘"High"‘, ‘"Medium"‘, or ‘"Low"‘), and optional ‘date_due‘ (datetime object).
@@
B.2 Task organization
Since each participant experienced both the baseline and proactive conditions, we distributed the 4 tasks in a symmetrical manner. This means that if the participant started with a system-building task (Task 1 or 2), they would see the other system-building task when the proactive assistant was switched on or off. This scheme ensures that we have a relatively similar number of each task type in both proactive and baseline conditions.
Appendix C Prompts
Debugging prompt
[insert code]
I need help identifying and fixing the issues in the code. Here are some relevant details:
stdout: [insert err.stdout]
stderr: [insert err.stderr]
Suggest 1 to 2 solutions to fix the issues.
Use the following format for your response: 1. short title explanation on what went wrong and how to fix it “‘ code snippet addressing the problem “‘
Preview prompt
We have the following ORIGINAL code:
[insert code]
A suggestion has been provided:
[insert suggestion]
INSTRUCTION: - If the suggestion contains no code or potential modifications to the code, return 0 and nothing else. Only zero - If the suggestions contain code, integrate the suggestion into the ORIGINAL code. Make minimal changes to the ORIGINAL code. Return the changed code only with nothing else.
Suggestion generation prompt
We use the following system prompt:
You are a helpful programming assistant. You can predict what the user wants to do next and provide suggestions. You never make mistakes in your code. Consider things such as explaining existing code, brainstorming new ideas, providing syntax hints and links to external documentations, identifying and fixing bugs, adding unit tests, completing unfinished code, improving code efficiency and modularity, etc. Do not include suggestions only adding docstrings. Do not include suggestions involving packages outside of vanilla python and numpy.
We use the following main prompt:
[insert code]
Provide suggestions based on the above code context. Include ‘function name‘ in suggestion title if it is mentioned in the suggestion.
Use one of the following formats depending on suggestion type, provide 3 suggestions:
1. suggestion type: short title explaining provided code or brainstorming high-level ideas
OR
1. suggestion type: short title “‘ one or more suggested code snippets, do not include ‘python‘ “‘ clear and detailed explanation for each code snippet in bullet point format. If code is very straightforward then don’t explain
"suggestion type" can be one of (Code Explanation, Code Improvement, Brainstorming Idea, Testing, Bug Fix, Syntax Hint) or something else. If there are multiple functions or classes in the code, provide references to the specific function or class that the suggestion pertains to in the explanation.

### v2

@@ 7.1.2. Expanding design considerations
(6) Allow users to decide when they want proactive assistance. Multiple participants across conditions recommended that they would like the option for when to have the proactive assistant on (e.g., “an opt-in feature” or “could change the frequency”), particularly because “sometimes I found it useful and sometimes I didn’t”. Since the goal of our study was to investigate the usage patterns and effects of a proactive assistant, providing the user the option to turn off the proactive assistant would not have been possible. This may be a particularly necessary feature in certain contexts where the user is working on less standard coding tasks, the user may choose to turn off proactive suggestions. Prior evaluations of LLMs have shown that models are more accurate or hallucinate less on tasks akin to those that have appeared in the training data (Jain et al., 2024; Liu et al., 2024a). This feature could also be important for long-term usage and facilitating appropriate trust and reliance on AI assistance. For example, a participant was concerned about potential over-reliance on proactive suggestions, which “can lead to a lack of understanding of a codebase. Therefore, I don’t want constant suggestions.”
@@
(7) Incorporate additional context when generating suggestions beyond user code. While many participants already appreciated the benefits of proactive suggestions, they had further ideas for improving the suggestion content by incorporating other aspects of user behavior. Incorporating additional context can not only better customize suggestions to user needs, but may also be beneficial in settings where the user is working on long context tasks. For example, participants suggested that the proactive assistant could take into account “the functions I was using’’, “what the user is focusing on, typing, or clicking,” or “recently updated code regions.” Another option suggested to incorporate further context would be to allow users to provide a natural language of their task—i.e., “if I can enter the task description somewhere as a constant part of the prompt [as] some suggestions have a slight misunderstanding of the problem”.
@@
(8) Flexibly varying the amount of information shown. Since the current proactive assistant provides one implementation option and provides a fixed number of suggestions each time, participants noted the possibility of scaling the amount of information up and down. For example, the assistant could provide more implementation options in the proactive suggestions, e.g., “multiple ways of solving a problem,” particularly when different factors may matter in different contexts (e.g., runtime). However, there may also be times when the assistant could scale down the number of suggestions shown (e.g., “if everything was good, then tell me that everything is good.”)
7.1.3. Comparison of findings to prior work
Finally, we discuss our findings in the context of related, prior literature on AI-assisted programming and reflect on how our results should implicate the design of future AI coding assistants. In particular, we focus on related work of two categories: findings on chat-based programming assistants and findings on other forms of proactive programming assistants.
Implications on chat-based programming assistants.
In a prior study on chat assistance, Ross et al. (2023) found that the “acceptance rate” of non-proactive chat requests, as measured by code copies, was higher than the acceptance rate of proactively generated code completions. However, our findings suggest that proactive chat assistants can generally increase user productivity. While these results may appear to be at odds, that is not necessarily the case as proactive suggestions can have a lower acceptance rate but still help users indirectly. Relatedly, Nam et al. (2024) and Mozannar et al. (2024c) both highlighted issues with and the burden of appropriately prompting the chat assistant to receive helpful responses. Our study provides evidence that a proactive assistant can reduce such a burden with proactive suggestions surfacing ideas that the user may not have thought of. In sum, our findings suggest that chat-based assistants may benefit from incorporating some aspects of proactivity.
Implications on proactive programming assistants.
Traditional approaches to providing proactive feedback relied on collecting relevant training examples and thus were largely limited in scope in terms of the types of feedback that the system could provide (Hartmann et al., 2010; Head et al., 2017). In contrast, we find that with modern LLMs, our proactive assistant could provide a diverse set of suggestions that were utilized to varying degrees by participants at different points of solving the programming task. This suggests the potential for leveraging current and even more capable models in the future as an engine for surfacing high-quality suggestions. Ferdowsi et al. (2024) showed previewing AI suggestions and their impacts on existing code facilitates user validation of these suggestions. While we do observe benefits from the Suggest and Preview condition, participants saw the biggest benefit from the Suggest condition where there was no preview functionality. A possible reason for this may be due to the way we present previews of code edits to users in our interface, which may not be optimal for user validation and is a ripe direction for future work.
7.2. Limitations and Future Work
@@
There are a few limitations to our study results and setup which we discuss here. First, our study was time-limited and consisted of tasks given to the participants rather than the tasks being self-motivated by the participants. The coding tasks we relied on do not span the entire set of tasks a professional programmer might encounter in their work. The tasks were also easily measured in the number of test cases passed, which may not be the best metric in practice. Second, the participants in our study were entirely comprised of undergraduate and graduate students, which is not fully representative of the audience that uses AI tools for programming. Third, our implementation of the proactive assistant was integrated into a web-based coding environment which does not contain the full features of an enterprise IDE. Fourth, the conditions that we considered in our controlled lab study were limited to chat-based systems, since not all participants had experience using GitHub Copilot. Altogether, we encourage the reader to consider the limitations of our study design when interpreting how our findings may generalize to various real-world programming applications. For example, while we observe that participants did not mention the correctness of suggestions, this may be an important consideration for developers who work on safety-critical code when interacting with a proactive assistant.
@@
Our study presents multiple opportunities for future work. First, further work is necessary to understand the nuances of proactive assistants on a wider range of user backgrounds and coding tasks. Additionally, future work could consider evaluating the proactive chat assistants in the presence of other programming tools like code completion to understand the interplay between different tools and their relative effects on productivity. There are also multiple promising directions in how proactive assistants can be improved to handle longer contexts, rank suggestions, more clever ways to trigger suggestions, and reduce potential hallucinations in assistant suggestions.
8. Conclusion
In this paper, we explore the design and implementation of proactive AI assistants powered by large language models for chat-based programming assistants. Our prototype considers a set of design considerations that balance the benefits of proactivity while handling its challenges. In a randomized experimental study, we evaluated the impact of various design elements of the proactive assistant on programmer productivity and user experience. We find that proactive assistants can generally increase the number of tasks completed and improve the user experience while coding. Furthermore, we provide insight into how aspects of the design including the effect of suggestion timing and the level of assistant involvement change interaction patterns. These results allow us to revisit the design implications for proactive coding assistants, validating the importance of the design considerations while introducing new considerations that allow for further customization and adaptivity to the user needs. Altogether, this work demonstrates the value of proactivity in AI coding assistants and identifies key directions for further study to further enhance user productivity and experience.
@@ A.1. Instructions
  • You can run your code by pressing the ”Run” button and the output will be in the output box at the bottom in grey.
@@
  • Pressing the ”Clear” button on the top of the chat window will clear the current chat history. Chat history will be automatically cleared when you move to the next task.
@@
  • Previewing: Click ”Preview” to see how the proactive assistant incorporates the suggestion into your code, which can then accept or hide the changes. Use the copy icon to copy the code to your clipboard.
@@
You can review this tutorial again by clicking on the “Show Instructions” button at the top of the page, then clicking ”Next”.
A.2. Post-study Questionnaire
@@ B.1. Tasks
You are an application developer working on the backend of a to-do list app. You are working with some existing code your colleague left off. Each task takes in a ‘description‘ (string), ‘category‘ (string), ‘priority‘ (string, ‘”High”‘, ‘”Medium”‘, or ‘”Low”‘), and optional ‘date_due‘ (datetime object).
@@
Appendix C Additional results
Figure 7. Percentage of sub-tasks completed correctly broken down by sub-groups. Comparing baseline chat to proactive assistants across different sub-groups as defined by Python programming experience, AI tool usage frequency, and gender. We report mean performance (by percentage sub-tasks completed) as well as standard error).
[…]
Percentage of sub-tasks completed correctly broken down by sub-groups. Comparing baseline chat to proactive assistants across different sub-groups as defined by Python programming experience, AI tool usage frequency, and gender. We report mean performance (by percentage sub-tasks completed as well as standard error).
Do the benefits of proactivity vary by user background?
As discussed in Section 6.1, we generally observe benefits across all proactive conditions. A natural follow-up is to analyze whether user benefits vary depending on user background. As part of background and demographic information, we collected participants’ self-reported Python proficiency, self-reported AI tool experience, and gender. For the purposes of this analysis, we combine all proactive conditions given the sample size per condition of each subgroup. We compute the difference in performance per participant between the proactive and baseline condition and build a linear model with the background and demographic information as fixed effects. We do not find any statistically significant effect due to Python experience or AI tool usage—this may be due to insufficient sample size per subgroup (e.g., there are only a small number of beginner programmers). However, we do observe a significant difference ($p=0.04$) due to gender where women tend to benefit more from proactive suggestions by 24.6%. This observation aligns with prior work that showed that software design could have different impacts on different genders (Burnett et al., 2016). In Figure 7, we visualize the percentage of sub-tasks completed correctly broken down by each of the sub-groups.
What kinds of suggestions do participants tend to accept?
As discussed in Section 6.3.3, of the 69 accepted suggestions, the most commonly types of accepted suggestions were those on brainstorming new functionality and debugging (latent errors), with 18 occurrences each. The least likely accepted suggestions were explaining existing code and pointers to documentation, with only 1 occurrence each. Other number of occurrences of other suggestions were 12 for completing unfinished code, 3 for debugging (runtime errors), 9 for adding unit tests, and 8 for improving efficiency and modularity. We also observe that all accepted suggestions contain code snippets.